# Some Geometric Properties of the Yff Points of a Triangle International Journal of Computer Discovered Mathematics (IJCDM) ISSN 2367-7775 ©IJCDM Volume 11, 2026 pp. xx–yy

### v1

@@ 1. Introduction
Theorem 1.1.
Let $D_{1}$, $D_{2}$, $E_{1}$, $E_{2}$, $F_{1}$, $F_{2}$ be points on the sides of triangle $ABC$ (as shown in Figure 1) such that $AE_{1}=BF_{1}=CD_{1}=AF_{2}=BD_{2}=CE_{2}=u$ where $u$ is the real root of the equation $x^{3}=(a-x)(b-x)(c-x)$. Then $AD_{1}$, $BE_{1}$, $CF_{1}$ meet in a point $Y_{1}$ and $AD_{2}$, $BE_{2}$, $CF_{2}$ meet in a point $Y_{2}$.
@@
Let $ABC$ be an arbitrary triangle (Figure 2). Then $Y_{1}Y_{2}\perp X_{1}X_{3}$.
@@ 2. Right Triangles
Let $ABC$ be a right triangle (named counterclockwise) with right angle at $B$ (Figure 3). Then $AY_{2}$, $CY_{1}$, and $X_{8}X_{20}$ are concurrent.
@@
Set up a barycentric coordinate system using $\triangle ABC$ as the reference triangle, so that he coordinates for $A$, $B$, and $C$ are $(1:0:0)$, $(0:1:0)$, and $(0:0:1)$, respectively. The coordinates for $X_{8}$ and $X_{20}$ can be found in [5].
@@
where $u$ is given in Theorem 1.1. We will call these the symmetric coordinates for $Y_{1}$ and $Y_{2}$.
@@
Theorem 2.2.
Let $ABC$ be a right triangle (named counterclockwise) with right angle at $B$ and $\angle ACB=30^{\circ}$ (Figure 4). Then $AY_{1}$, $BY_{2}$, and $X_{8}X_{21}$ are concurrent.
@@
Let $ABC$ be a right triangle with right angle at $B$ (Figure 5). Then $BX_{8}\perp Y_{1}Y_{2}$.
@@ 3. $60^{\circ}$ Triangles
Let $ABC$ be a triangle (named counterclockwise) with $\angle C=60^{\circ}$ (Figure 6). Then $AY_{1}$, $BY_{2}$, and $X_{3}X_{8}$ are concurrent.
@@ 4. Heptagonal Triangles
Let $ABC$ be a heptagonal triangle with $a>b>c$. Then
@@
Let $P$ have barycentric coordinates $(p:q:r)$ with respect to $\triangle ABC$. Then $Q$, the reflection of $P$ about side $BC$, has coordinates
@@
that is, the infinite point of $PX$ is the infinite point of the AA-altitude and then $PX$ is perpendicular to $BC$.
@@
Let $ABC$ be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 7). Let $Y_{1}^{\prime}$ be the reflection of $Y_{1}$ about side $BC$. Then $X_{1}Y_{2}\parallel Y_{1}^{\prime}C$.
@@
Theorem 4.4.
Let $ABC$ be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 8). Then $\angle CX_{1}Y_{2}+\angle X_{1}CY_{1}=\frac{\pi}{7}$.
@@
Let $Y_{1}^{\prime}$ be the reflection of $Y_{1}$ about $BC$. Draw in line $CX_{3}$ as shown in Figure 9.
@@
Since $X_{3}$ is the center of the circle through $A$, $B$, and $C$, $\angle CX_{3}B=6\pi/7$. This implies $\angle BCX_{3}=\pi/14$. Since $X_{1}$ is the incenter, $CX_{1}$ bisects $\angle ACB$, so $\angle X_{1}CB=\pi/14$. Therefore $\angle BCX_{3}=\angle X_{1}CB$.
Since $Y_{1}^{\prime}$ is the reflection of $Y_{1}$ about side $BC$, $\angle Y_{1}CB=\angle BCY_{1}^{\prime}$, i.e. the dotted angles are equal (the “black angles”). Subtracting, we find that $\angle X_{1}CY_{1}=\angle Y_{1}^{\prime}CX_{3}$, i.e. the green angles are equal.
@@
Let $ABC$ be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$. Let the angles be named as shown in Figure 10. If $\angle Y_{2}X_{1}A=\theta$, then
@@
Let $ABC$ be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 11). Then $\angle CX_{1}Y_{2}=\angle Y_{1}CX_{3}$.
@@
Since $X_{3}$ is the center of the circle through $A$, $B$, and $C$, $\angle CX_{3}B=6\pi/7$. This implies $\angle BCX_{3}=\pi/14$. Then, using Theorem 4.5, we have $\angle Y_{1}CX_{3}=\angle BCX_{3}+\angle 6=\pi/14+4\pi/7-\theta=9\pi/14-\theta=\angle 2$. ∎
@@
Let $ABC$ be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$. Let $P$ be the foot of the perpendicular from $Y_{1}$ to $AC$ (Figure 12). Then $\angle CBX_{5}=\angle Y_{2}Y_{1}P$.
@@
Let $ABC$ be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 13). Then $\angle X_{21}X_{1}Y_{2}=\angle CY_{1}Y_{2}$.
@@
Let $ABC$ be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 14). Then $\angle X_{1}Y_{2}Y_{1}=\angle X_{28}CY_{1}$.
@@ 5. Harmonic Triangles
Let $ABC$ be a triangle such that $Y_{1}Y_{2}$ passes through a vertex. Then $Y_{1}Y_{2}$ is a median of the triangle.
@@
Let $BY_{1}$ meet $AC$ at $M$. Then, from Figure 1 (left), we have $CM=u$. Line $BY_{2}$ is the same line, so also meets ACAC at $M$. From Figure 1 (right), we have $AM=u$. Thus, $AM=CM$ and $BM$ is a median of the triangle. ∎
@@
Let $ABC$ be a triangle such that $BY_{1}$ is a median. Then $Y_{1}$ divides the median into two segments in the ratio 2b:2ab2b:2a-b.
@@
Let $BY_{1}$ meet $AC$ at $E$ and let $AY_{1}$ meet $BC$ at $D$.
Then, from Figure 1 we see that $CE=BD=u$. Since $BY_{1}$ is a median, $AE=CE$, so $AE=u$. Since $BD=u$, we have CD=auCD=a-u. Line $AD$ is a transversal of $\triangle EBC$, so by the Theorem of Menelaus, we have x(au)u=yu(2u)x(a-u)u=yu(2u), so x/y=2u/(au)x/y=2u/(a-u). But $u=b/2$. Thus $x/y=b/(a-\frac{b}{2})$ which implies x/y=2b/(2ab)x/y=2b/(2a-b). ∎
@@
Let $ABC$ be a triangle such that $BY_{2}$ is a median. Then $Y_{2}$ divides the median into two segments in the ratio 2(2ab):b2(2a-b):b.
@@
Without loss of generality. assume that $Y_{1}Y_{2}$ passes through $E$, the midpoint of $AC$. The barycentric coordinates for $E$ are $(1:0:1)$. Using the simple coordinates for $Y_{1}$ and $Y_{2}$, the determinant condition for $Y_{1}$, $Y_{2}$ and $E$ to be collinear is
@@ 7. Double-Angle Triangles
Let $ABC$ be a triangle in which $\angle A=2\angle B$. Then $a^{2}=b(b+c)$.
@@
Let $\angle CBA=\theta$, so that $\angle BAC=2\theta$. Extend $CA$ past $A$ to a point $D$ such that $AD=AB$, so that $ABD$ is an isosceles triangle and $\angle ABD=\angle BDA$.
@@
Let $ABC$ be a double-angle triangle (named counterclockwise) with $\angle A=2\angle B$ (Figure 17). Let $Y_{2}^{\prime}$ be the reflection of $Y_{2}$ about side $BC$. Then $BY_{2}$, $CX_{2}$, and $AY_{2}^{\prime}$ are concurrent.
@@ 8. Isosceles Triangles
Let $Y$ be one of the Yff points of $\triangle ABC$. Let the areas of triangles $BCY$, CAYCAY, and $ABY$ be $K_{a}$, $K_{b}$, and $K_{c}$, respectively. Then $K_{b}K_{c}=K_{a}^{2}$ if and only if $\triangle ABC$ is isosceles with $AB=AC$.
@@ 9. Isosceles Right Triangles
Let $ABC$ be an isosceles right triangle (named counterclockwise) with right angle at $A$. Let $X_{1}Y_{2}$ meet $BC$ at $D$ (Figure 18). Then $\angle DY_{2}C=\angle CBX_{8}$.
@@
Let $ABC$ be an isosceles right triangle (named counterclockwise) with right angle at $A$. Then $\angle Y_{2}X_{1}C=\angle Y_{2}CX_{9}$ (Figure 19).
@@ 10. Special Triangles
Let $BX_{8}$ meet $AC$ at $E$. Then by a well-known property of the Nagel point, CE=saCE=s-a where $s$ is the semiperimter of the triangle [9, §3.2.2].
Then $Y_{1}$ lies on $BE$, if and only if sa=us-a=u. Eliminating $u$ between this condition and Equation (1) gives the condition to be $a^{2}(a-b-c)=bc(b+c-3a)$ as required. ∎
@@ 11. Cevian Triangles
In $\triangle ABC$, let $D$ be a point on $BC$ and let $E$ be a point on $AC$ (Figure 21). Then
@@
Let $P$ be a point inside $\triangle ABC$. Let $AP$ meet $BC$ at $D$, BPBP meet $CA$ at $E$ and $CP$ meet $AB$ at $F$. Suppose $[ABC]=K$, $AB=c$, $BC=a$, $CA=b$, $BD=a_{1}$, $DC=a_{2}$, $CE=b_{1}$, $EA=b_{2}$, $AF=c_{1}$, and $FB=c_{2}$ as shown in Figure 22. Then
@@
Triangle $DEF$ is known as the cevian triangle of point $P$.
@@
In $\triangle ABC$, let $AY_{1}$ meet $BC$ at $D$, $BY_{1}$ meet $CA$ at $E$ and $CY_{1}$ meet $AB$ at $F$ (Figure 23). Then
@@ 12. Special Points
Let $Y_{c}$ be the center of the unique conic that passes through the points $A$, $B$, $C$, $Y_{1}$, and $Y_{2}$. The point $Y_{c}$ is not a triangle center listed in ETC [5] as of January 1, 2026. The first normalized trilinear coordinate of $Y_{c}$ in a 6–9–13 triangle is 0.7440584232553069.

### v2

@@ 1. Introduction
Theorem $1.$
Let $D_{1}$, $D_{2}$, $E_{1}$, $E_{2}$, $F_{1}$, $F_{2}$ be points on the sides of triangle ABCABC (as shown in Figure 1) such that $AF_{1}=BD_{1}=CE_{1}=AE_{2}=BF_{2}=CD_{2}=u$ where $u$ is the real root of the equation $x^{3}=(a-x)(b-x)(c-x)$. Then $AD_{1}$, $BE_{1}$, $CF_{1}$ meet in a point $Y_{1}$ and $AD_{2}$, $BE_{2}$, $CF_{2}$ meet in a point $Y_{2}$.
@@
Let ABCABC be an arbitrary triangle (Figure 2). Then $Y_{1}Y_{2}\perp X_{1}X_{3}$.
@@ 2. Right Triangles
Let ABCABC be a right triangle (named counterclockwise) with right angle at $B$ (Figure 3). Then $AY_{2}$, $CY_{1}$, and $X_{8}X_{20}$ are concurrent.
@@
Set up a barycentric coordinate system using $\triangle ABC$ as the reference triangle, so that he coordinates for AA, $B$, and $C$ are $(1:0:0)$, $(0:1:0)$, and $(0:0:1)$, respectively. The coordinates for $X_{8}$ and $X_{20}$ can be found in [5].
@@
where $u$ is given in Theorem $1.$ We will call these the symmetric coordinates for $Y_{1}$ and $Y_{2}$.
@@
Theorem $2.$
Let ABCABC be a right triangle (named counterclockwise) with right angle at $B$ and $\angle ACB=30^{\circ}$ (Figure 4). Then $AY_{1}$, $BY_{2}$, and $X_{8}X_{21}$ are concurrent.
@@
Let ABCABC be a right triangle with right angle at $B$ (Figure 5). Then $BX_{8}\perp Y_{1}Y_{2}$.
@@ 3. $60^{\circ}$ Triangles
Let ABCABC be a triangle (named counterclockwise) with $\angle C=60^{\circ}$ (Figure 6). Then $AY_{1}$, $BY_{2}$, and $X_{3}X_{8}$ are concurrent.
@@ 4. Heptagonal Triangles
Let ABCABC be a heptagonal triangle with $a>b>c$. Then
@@
Let $P$ have barycentric coordinates $(p:q:r)$ with respect to $\triangle ABC$. Then $Q$, the reflection of $P$ about side BCBC, has coordinates
@@
that is, the infinite point of PXPX is the infinite point of the AA-altitude and then PXPX is perpendicular to BCBC.
@@
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 7). Let $Y_{1}^{\prime}$ be the reflection of $Y_{1}$ about side BCBC. Then $X_{1}Y_{2}\parallel Y_{1}^{\prime}C$.
@@
Theorem $4.$
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 8). Then $\angle CX_{1}Y_{2}+\angle X_{1}CY_{1}=\frac{\pi}{7}$.
@@
Let $Y_{1}^{\prime}$ be the reflection of $Y_{1}$ about BCBC. Draw in line $CX_{3}$ as shown in Figure 9.
@@
Since $X_{3}$ is the center of the circle through AA, $B$, and $C$, $\angle CX_{3}B=6\pi/7$. This implies $\angle BCX_{3}=\pi/14$. Since $X_{1}$ is the incenter, $CX_{1}$ bisects $\angle ACB$, so $\angle X_{1}CB=\pi/14$. Therefore $\angle BCX_{3}=\angle X_{1}CB$.
Since $Y_{1}^{\prime}$ is the reflection of $Y_{1}$ about side BCBC, $\angle Y_{1}CB=\angle BCY_{1}^{\prime}$, i.e. the dotted angles are equal (the “black angles”). Subtracting, we find that $\angle X_{1}CY_{1}=\angle Y_{1}^{\prime}CX_{3}$, i.e. the green angles are equal.
@@
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$. Let the angles be named as shown in Figure 10. If $\angle Y_{2}X_{1}A=\theta$, then
@@
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 11). Then $\angle CX_{1}Y_{2}=\angle Y_{1}CX_{3}$.
@@
Since $X_{3}$ is the center of the circle through AA, $B$, and $C$, $\angle CX_{3}B=6\pi/7$. This implies $\angle BCX_{3}=\pi/14$. Then, using Theorem 4.5, we have $\angle Y_{1}CX_{3}=\angle BCX_{3}+\angle 6=\pi/14+4\pi/7-\theta=9\pi/14-\theta=\angle 2$. ∎
@@
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$. Let $P$ be the foot of the perpendicular from $Y_{1}$ to ACAC (Figure 12). Then $\angle CBX_{5}=\angle Y_{2}Y_{1}P$.
@@
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 13). Then $\angle X_{21}X_{1}Y_{2}=\angle CY_{1}Y_{2}$.
@@
Let ABCABC be a heptagonal triangle (named counterclockwise) with $AB<AC<BC$ (Figure 14). Then $\angle X_{1}Y_{2}Y_{1}=\angle X_{28}CY_{1}$.
@@ 5. Harmonic Triangles
Let ABCABC be a triangle such that $Y_{1}Y_{2}$ passes through a vertex. Then $Y_{1}Y_{2}$ is a median of the triangle.
@@
Let $BY_{1}$ meet ACAC at $M$. Then, from Figure 1 (left), we have $CM=u$. Line $BY_{2}$ is the same line, so also meets ACAC at $M$. From Figure 1 (right), we have $AM=u$. Thus, $AM=CM$ and BMBM is a median of the triangle. ∎
@@
Let ABCABC be a triangle such that $BY_{1}$ is a median. Then $Y_{1}$ divides the median into two segments in the ratio 2b:2ab2b:2a-b.
@@
Let $BY_{1}$ meet ACAC at $E$ and let $AY_{1}$ meet BCBC at $D$.
Then, from Figure 1 we see that $CE=BD=u$. Since $BY_{1}$ is a median, $AE=CE$, so $AE=u$. Since $BD=u$, we have CD=auCD=a-u. Line ADAD is a transversal of $\triangle EBC$, so by the Theorem of Menelaus, we have x(au)u=yu(2u)x(a-u)u=yu(2u), so x/y=2u/(au)x/y=2u/(a-u). But $u=b/2$. Thus $x/y=b/(a-\frac{b}{2})$ which implies x/y=2b/(2ab)x/y=2b/(2a-b). ∎
@@
Let ABCABC be a triangle such that $BY_{2}$ is a median. Then $Y_{2}$ divides the median into two segments in the ratio 2(2ab):b2(2a-b):b.
@@
Without loss of generality. assume that $Y_{1}Y_{2}$ passes through $E$, the midpoint of ACAC. The barycentric coordinates for $E$ are $(1:0:1)$. Using the simple coordinates for $Y_{1}$ and $Y_{2}$, the determinant condition for $Y_{1}$, $Y_{2}$ and $E$ to be collinear is
@@ 7. Double-Angle Triangles
Let ABCABC be a triangle in which $\angle A=2\angle B$. Then $a^{2}=b(b+c)$.
@@
Let $\angle CBA=\theta$, so that $\angle BAC=2\theta$. Extend CACA past AA to a point $D$ such that $AD=AB$, so that ABDABD is an isosceles triangle and $\angle ABD=\angle BDA$.
@@
Let ABCABC be a double-angle triangle (named counterclockwise) with $\angle A=2\angle B$ (Figure 17). Let $Y_{2}^{\prime}$ be the reflection of $Y_{2}$ about side BCBC. Then $BY_{2}$, $CX_{2}$, and $AY_{2}^{\prime}$ are concurrent.
@@ 8. Isosceles Triangles
Let $Y$ be one of the Yff points of $\triangle ABC$. Let the areas of triangles BCYBCY, CAYCAY, and ABYABY be $K_{a}$, $K_{b}$, and $K_{c}$, respectively. Then $K_{b}K_{c}=K_{a}^{2}$ if and only if $\triangle ABC$ is isosceles with $AB=AC$.
@@ 9. Isosceles Right Triangles
Let ABCABC be an isosceles right triangle (named counterclockwise) with right angle at AA. Let $X_{1}Y_{2}$ meet BCBC at $D$ (Figure 18). Then $\angle DY_{2}C=\angle CBX_{8}$.
@@
Let ABCABC be an isosceles right triangle (named counterclockwise) with right angle at AA. Then $\angle Y_{2}X_{1}C=\angle Y_{2}CX_{9}$ (Figure 19).
@@ 10. Special Triangles
Let $BX_{8}$ meet ACAC at $E$. Then by a well-known property of the Nagel point, CE=saCE=s-a where $s$ is the semiperimter of the triangle [9, §3.2.2].
Then $Y_{1}$ lies on BEBE, if and only if sa=us-a=u. Eliminating $u$ between this condition and Equation (1) gives the condition to be $a^{2}(a-b-c)=bc(b+c-3a)$ as required. ∎
@@ 11. Cevian Triangles
In $\triangle ABC$, let $D$ be a point on BCBC and let $E$ be a point on ACAC (Figure 21). Then
@@
Let $P$ be a point inside $\triangle ABC$. Let APAP meet BCBC at $D$, BPBP meet CACA at $E$ and CPCP meet ABAB at $F$. Suppose $[ABC]=K$, $AB=c$, $BC=a$, $CA=b$, $BD=a_{1}$, $DC=a_{2}$, $CE=b_{1}$, $EA=b_{2}$, $AF=c_{1}$, and $FB=c_{2}$ as shown in Figure 22. Then
@@
Triangle DEFDEF is known as the cevian triangle of point $P$.
@@
In $\triangle ABC$, let $AY_{1}$ meet BCBC at $D$, $BY_{1}$ meet CACA at $E$ and $CY_{1}$ meet ABAB at $F$ (Figure 23). Then
@@ 12. Special Points
Let $Y_{c}$ be the center of the unique conic that passes through the points AA, $B$, $C$, $Y_{1}$, and $Y_{2}$. The point $Y_{c}$ is not a triangle center listed in ETC [5] as of January 1, 2026. The first normalized trilinear coordinate of $Y_{c}$ in a 6–9–13 triangle is 0.7440584232553069.